Construct the confidence interval for the population mean μ cz 0.90, x-9.6, σ 0.4, and n 51 A 90% confidence interval for μ is (LL) (Round to two decimal places as needed.)
Answer #1

The 90% confidence interval for population mean is given by:

x ±20.95 × Vn

Now, z​​​​​​0.95 is 1.645

Hence, the 90% confidence interval is

9.6 1.645 × 9.69 0.4 -= /51

And

0.4 9.6-1.645 × = 9.51 /51

Hence, a 90% confidence interval for mu is (9.51,9.69)